A 1511 kg car is traveling down the road at 81.0 km/h. While traveling at this rate of speed, what is the kinetic energy of this vehicle in kilojoules?
1
Expert's answer
2019-03-07T07:04:24-0500

KE = 0.5mv^2 = 0.5 * 1511kg * (81 kmph)^2

KE = 0.5 * 1511kg * 81^2 * (1000m / 3600s)^2 = 382.5 kJ